FEL Upgrade February 2002 Monthly Report

His month the optics team demonstrated that our proposed scheme for radius of curvature control of the primary optical cavity mirrors for the 10 kW Upgrade works extremely well in terms of resolution, linearity and concentricity. Details are given in the Optics Section below. The procurement package for the optical cavities for the 10 kW Upgrade were sent out this month.

First estimates for the 40 K helium flow required to cool the cryogenic sapphire outcoupler were encouraging, < 1 gm/sec. This is suggests it will be relatively easy to cool not only the 1 micron IR system but also the UV and scaled up IR lasers using this approach.

Significant progress on the gun occurred with the chamber going out for vacuum firing and the support tube getting the inner ring welded.

The first production coil for the optical chicane dipoles was successfully produced.

Magnet stand position data was delivered to the alignment team so they can establish installation locations in the tunnel in the next couple of weeks. Stand hardware has been fabricated and is ready to install in many places so this effort can get underway when the survey is complete.

Beam diagnostics is making good progress with component assembly: sixteen viewing windows and six assembled shielded viewers have been delivered for assembly onto girders.

We completed revisions of the our primary contractual agreement between the Navy and DOE for the FEL activities at Jefferson Lab (the "MOA"), that would fund additional work in FY02 and FY03 including:

(1) commissioning of the FEL Upgrade
(2) design and construction of a short wavelength (1-2 micron) wiggler
(3) completion of the 3rd cryomodule for the FEL linac
(4) upgraded drive laser for the photocathode gun (JTO funded)
(5) design and testing of cryogenically cooled optical cavity mirrors (JTO funded)

The revised MOA includes work statements and cost summaries for the above activities. These documents have been forwarded to our DOE and ONR contracting officers for approval. A draft rebaselined Phase II budget was produced for Program Manager review which incorporated changes

consistent with the funding available in the revised MOA.

Efforts are underway to incorporate each WBS element's installation schedule as based on expected delivery and checkout of hardware into a master installation plan. Also to be incorporated are outside activities that relate to resource loading such as CEBAF operation schedule, etc.

Discussions were held on approaches for wiggler replacement for 1 micron output optimization. UV cavity approaches were the subject of a separate meeting. Analysis and discussions in both of those areas continues.

FEL project management and AES, Inc have agreed upon the joint work statement for a CRADA on 100 mA injector development. The document goes on for the final packaging needed for the governmental (DOE/DOD) reviews. A support letter enumerating the JLab tasks was forwarded to AES.

For the high power injector work, proposed 750 MHz injector cavity shapes were received from AES for review and analysis for multipacting.

This is the last week of the 2002 session of the Virginia General Assembly, so we had to respond to several requests for information concerning our FEL activities that are funded by the Commonwealth of Virginia.

A visit was made to Northrop Grumman-Newport News this month to discuss potential collaboration on the proposed 100 kW Upgrade.

We had a visit from representatives from Gas Technology Institute who were interested in possible applications of the FEL for gas and oil extraction.

F. Dylla visited the Wellman Lab of Photomedicine on Feb. 13 to present a seminar and discuss potential collaborations on FEL applications to bio-medicine. He visited the Engineering School at Northeastern University on Feb.15 to present a seminar on materials processing with the FEL.

The Program Announcement is out for the fall Directed Energy Professional Society in Monterey November 12-15 and we are encouraging participation of the program team. There will be a day course on FELs as well as a good session to include our JTO and Navy efforts.

Project Cost Performance:
The Phase 1 FEL Upgrade project budget for the period June 1, 2000 to September 30, 2001 is $9,029k. The first phase of the project was fully obligated as of the end of September. The total Phase 1 project cost to date-actual costs plus commitments is $9,029k as planned for Phase 1 completion.

The Phase 2 FEL Upgrade project was approved for $4,500k for a performance period of February 1, 2001 to September 30, 2002. Actual funds transferred from ONR to DOE for this effort in FY02 were $4,442k. Work for Phase 2 proceeded to plan during this reporting period. Actual charges of $288k were accrued for February for total accrued costs of $2364k to date for Phase 2. The bulk of this effort involves procurements of components designed during Phase 1 for delivery and installation in Phase 2. Work scheduled for February was $327k for a total of $1019k for Phase 2 to date; work performed for February was $439k for a total of $1234k for Phase 2 to date, which results in 23% scheduled vs. 28% performed for the Phase 2 project. At present the cost variance for Phase 2 is -$1130k and the schedule variance is $215k. We expect to meet the planned cost at completion of Phase 2 effort. The program is presently on track to begin commissioning efforts in September 2002.

The UV FEL project was approved for $2,836k for a performance period from September 2001 to September 30, 2002. Actual charges of $168k were accrued through February for a total accrued costs of $233k since project start. Work scheduled for February was $276k for a total of $537k for the UV project to date, work performed for February was $227k for a total of $503k for the UV project to date, which results in 19% scheduled vs. 18% performed for the UV project. At present the cost variance for the UV project is $270k and the schedule variance is -$34k. The program is on track to meet the statement of work requirements on schedule and within the available budget.

WBS 3 (Beam Physics):
Work to clarify diagnostic and correction system configurations continues. In support of this, and commissioning plans, the Demo machine model is being upgraded to describe, well what else – the Upgrade.

Work continued on machine modeling, commissioning plan, and evaluation of magnet field quality (both simulated [dipoles] and measured [quadrupoles])

Consideration of endloop dipole (GX, GQ, GY) performance continues. All appear to have adequate field uniformity over the full operating range (80-210 MeV/c). All appear, however, to run somewhat saturated at full (210 MeV/c) excitation. As field and field integral matching was done at this operating point, all fail to stay within tracking specification (1 ppt) for field and field integral as excitations are lowered into the 80-145 MeV range.

Initial analysis suggests that GX & GQ magnets can be well matched in field integral for IR operation over 80-190 MeV/c by appropriate reduction of the steel length (about 3 mm for the GX and 5 mm for the GQ). This leads to very good field performance in the GQ (less than 1 ppt from design), while leaving the GX a factor of 2 or so out of spec in field. This is likely acceptable – the GX is at a location with modest beam envelopes and low dispersion; the focussing errors associated with deviations in field will not strongly influence machine performance. With "shaved" lengths at 210 MeV/c, field integrals in the GX and GQ fail to track at the 1/4% level, and core fields fail to track at the 1/3% level. Should this prove operationally unacceptable (either through excessive correction requirements or excessive focussing errors), shimming the GX with a single 2.5 mm shim (on the IR path, at the exit of the magnet) and shimming the GQ with a pair of 2.5 mm shims will bring both back within specification.

UV operation of the GX is also influenced by saturation. Core field matching at full excitation (210 MeV/c) will require ~2% offset of the bypass GX from the buss current; this falls to ~0.1% at 145 MeV/c. In either case, the magnet also fails to track in field integral by an amount consistent with ~3 mm length error. Should the aforementioned "shave" be done, this will be compensated at all excitations. Core field matching can in principle be accomplished through the use of a shunt. In practice, it may be less risky to energize the "shorted" coil using a trim supply, inasmuch as the required correction (0.1%) at low energy suggests that the shunt may run out of range and a bipolar trim be required if the magnetic model proves even slightly inaccurate.

GY performance poses a similar challenge - nominally designed to shunt 1.5 % of the current at full excitation, the required GY shunt falls to order 0.1% at 80 to 145 MeV/c. In practice, the magnet may run outside the available shunt range. This may be addressed two ways - a correction coil winding on the yoke and/or a small reduction in gap. Given that a single "backleg" turn will require 7m of wire, the latter solution may be preferable.

GW performance will be similarly examined to ensure it matches adequately with all this. George B. points out that not only will all this likely produce a properly functioning machine, it gives the WBS 3 manager magnets that are just far enough from spec that he can whine about them indefinitely, even though they'll work just fine.

WBS 4 (Injector):
Fit up of the internal parts for the ball cathode is being done prior to polishing. A mount for implanting the ball is being designed. The cathode support tube had the inner ring welded in and is having the oxide regions polished out. Part fit up and fine tuning continues. The ceramic stack had the replacement ceramic welded on, waiting on a leak check. The gun chamber was vacuum fired and the flanges cleaned; flange welding will complete the chamber. Continued work on the inline resistor design. The extractor cart drawings were signed off and submitted for bids.

PARMELA simulations of the transverse and longitudinal emittance evolution have been performed for 135 pC and 500 kV. So far the simulation has been run from the gun up to IMS0F03. The effect of the number of particles used for the simulations is under study. Similar simulations of the bunch length (in ps and mm) have been performed from the GUN up to MDU0F03 to provide information for ultra-fast laser imaging of the electron bunch. Efforts to implement PARMELA on PC continue.

A GaAs wafer was H2 cleaned for use in a field emission study. This is of interest for our quantification of cathode dark currents as we raise the gun voltage from 350kV in the 1 kW Demo to 500 kV in the Upgrade.

Previously taken data of an implanted electrode has been fitted with field emission equation. Results of the simulation show that the field enhancement factor beta ~ 260, the emitter height ~ 300 nm, emitter tip radius ~ 1 nm and the work function (which is unknown for the implanted coating) is very similar to that of stainless steel ~4.25 eV.

On the cleaned GaAs wafer, no signs of dust were observed in the SEM, although some "flake like" features appear scattered all over the surface. No emitters were found by the DC Field Emission Scan at 35 MV/m, except for a small emitter on the edge of the wafer. These preliminary results are encouraging since the electric field in the upgrade gun will be 6-10 MV/m. A new Scan is in progress at 60 MV/m.

Gun HVPS - The design of the combined Running & Conditioning Resistor mechanism is nearly finished. A design review will be held next week. The breakdown voltage and expected safety margin at several locations has been calculated. A cable tray, desk, table, and rack were moved out of the way in preparation for removing the old Gun HVPS Tank. The bids for the new HVPS Tank closed this month and a contract was placed with Industrial Alloy Fabrication located in Richmond. It is due May 10th.

WBS 6 (RF):
Quarter HVPS - Except for the HV capacitors and the PLC expansion units, all major components have been installed in each HVPS. The schematics are at the 98% level and the wiring has started.

Quarter Klystrons - The design for the klystron water manifold is finished and drawings are being updated. Two additional klystron carts will be constructed as soon as the drawings are finished. The drawings for the klystron carts are nearly complete. The old circulators and their controllers were removed this week. The new ones are to be installed next week. The correct flow meters for the RF Circulator and Load were ordered.

Injector HVPS - The Electricians finished wiring the high current primary side of the HVPS. New wiring was pulled and wired to the build's switchgear. For safety reasons, the final connection to the HVPS will not be made until it is needed to checkout the HVPS. Design of the PLC system and code is progressing.

Injector Klystrons - Brackets were made and used to install the new Circulators. Discussions are ongoing on how to install the larger LCW taps. New 8" LCW piping is being considered for the first 30' in the RF Gallery. The removal of an orifice to measure LCW flow is being considered also. These changes are to provide more flow for the upgrade equipment.

WBS 8 (Instrumentation):
Cable removal in the vault continues. We expect to have all of the cables we plan to remove out by the end of February and have begun plans to map magnet trim cables and power supply cables. Installation to begin around March 1.

The Machine Protection System input card design is complete and in to EECAD for final documentation prior to fabrication.

The technical requirements to shift the current 4 channel Beam Loss Monitoring system from a CAMAC to a VME based system are being determined and a plan is in progress to pursue this change.

New security features have been implemented to the FEL informational database. It can be accessed at: http://laser.jlab.org/data. Users with a JLAB e-mail address are permitted to log on immediately. Those users that do not have a JLAB e-mail address will be directed on the actions to take to become registered users of the system.

The drawings are being reviewed for the VME input card and fabrication is expected to begin soon. Drawings are also being revised for new beamviewer camera lamp rings to provide more light to the camera.

The machine shop has developed a device to put registration mark the foils on the beamviewers so when it is being used we will be better able to determine spot size and location on the viewer. New cameras are being built and the existing ones tested to support re-using them on the new machine.

Stepper motor control work continues. The chassis design is complete and fabrication is about to begin. The associated software is being developed.

A test fixture for the new CCD cameras has been built so each camera can be tested as it is built and will be ready for installation. A written procedure has been developed that outlines both camera assembly and testing.

Parts are on order for the new 8-channel OMS stepper motor chassis with fabrication expected to begin next month.

Initial mapping of the magnet power supplies is underway. Trim channels are also being designated as required. Cable routing for these items will be starting soon.

The MPS VME interface card check print is complete and design work continues in switching the Beam Loss Monitors from CAMAC to VME based.

The requirements and designs for the vacuum control system for the upgrade are well underway. Cable layouts and channel designations are complete. The chassis design is in the layout phase and should begin fabrication in April.

Sixteen viewing windows and six assembled shielded viewers have been delivered for assembly onto girders. Assembly of beamviewer control boxes is continuing and the boxes attached to the viewers as they are completed.

Nine additional beamviewers have been delivered for assembly onto girders bringing the total to fourteen.

The Beamloss Monitoring System design review is in progress. Procurement and fabrication is expected to begin soon. We are also moving into the design phase of an upgrade to the Analog Monitoring System. Work on a system drawing for the Laser Safety System has resumed as well.

Trim magnet cable mapping is well underway. We expect to begin pulling new cables and re-routing old ones as soon as the underway portion of the mapping of the floor in the vault is complete.

WBS 9 (Transport)
Optical Chicane Dipoles (GW)
• At Magnet Enterprises International in Oakland CA. the second coil was wound and is ready for potting early in the next month. The third coil is wound.
• Substantial progress has been mad in manufacturing the cores and parts using subcontractors who make similar parts for Stanford Linear Accelerator Center. Return legs are being ground to their final size. Top and bottom slabs will finish rough machining early next week. Field clamp material is in and those parts are being machined in house.
Injector Dipoles (DU/DV)
• Contracts were let to Wang NMR for coils and Assembly and to Master Machine for Cores and Parts for the Injector Dipoles.
• Master Machine will pick up material from us to start their manufacture. .
• DULY Research completed the Revision A of the drawings of both Injector Dipole (GU & GV) to give the final adjustment to meet specifications. On the GV, he height of return legs will need to change slightly and a small piece of mu metal is needed over the beam path adjacent to the horizontal field clamp parts. Production may start because the final machining of those parts that incorporates these changes is many weeks away. We are checking the changes (which do not preclude start of manufacturing) and plan to issue them the contractors the first week of March.
Arc Dipoles (GY, GX, GQ)
• Electronic procurement of Bend (GX) and Reverse Bend (GQ) Dipoles started. We gave vendors an additional week to respond because two vendors requested the extension. Bids are due the second week of March
• Advanced Energy Systems (AES) completed their analysis of all three magnets at low excitation for low energy for UV as well as IR.
• David Douglas almost completed his analysis of the use of the Bend and Reverse Bend Dipoles (GX & GQ) at all levels of excitation. His conclusion is that the faces may need to be shaved by about 2 mm with shimming back only necessary for use at 210 MeV/c.
• Wang NMR of CA manufactured the first GG Trom Coil and are now starting to make the GY coil winding mandrel. They have spooled and insulated all the conductor into 4-in-hand bundles so that winding may begin as soon as the form is available.
• Bosma Machine torch cut out all the core pieces, completed rough machining of the core pieces and sent them for heat treatment. Shim pieces are cut to size.
3 inch quad (QX)
• We concentrated on getting the magnet measurement stand qualified to measure the quadrupoles. We were confronted with data that has scatter of the same order as that of the tolerance in the specification for field quality. We learned how to measure the magnets and handle the data to reduce the scatter.
• Mechanical measurements show the machined contact surfaces are tied to the parting plane notches within a maximum of half of the one radian roll specification. These parting plane surfaces are within .001 inch of the pole tip edges. This indicates that we will be able to reliably use these surfaces for accurate alignment.
• We received the majority of the QX quadrupoles needed for the IR machine. The magnet test stand was qualified after we worked out our initial bugs. Eight quadrupoles were measured. All met the field quality requirements. However, they are a factor of 4 out of specification in matching from magnet to magnet that would allow one excitation curve to characterize them all. We worked it out with the software folks that it is OK to assign each quadrupole an individual excitation curves.
Trim Quad (QT)
• Milhous Control of Virginia started winding the first magnet. The custom magnet wire we supplied was out of specification because half of it was twisted beyond the ability to use it and the remainder was beyond the thickness specification. New wire we supplied restarted the process for the side coils but is too small for the top/bottom coils. The wire manufacturer will re-spool the twisted wire that is within size specification to remove the twist.
Sextupole (SF)
• DULY has a magnetic model that is within specification but has poles that are less wide that is ideal for fabrication and saturation.
• They are poised to do the analysis to produce the final design. They will lengthen the pole to reduce the saturation effects, widen it to match the pole face, reduce the coil turns to match the lower field and will optimize the sextupole integral by changing the exponent of the pole’s defining curve.
• By month’s end, they have not advanced because their analysis expert was back in Russia tending to his wife and newborn daughter.
Octupole (OT)
• Work on this magnet is on furlough until the sextupole is designed.
Beam Line and Vacuum
• We are working of the regions before the first arc and after the second arc.
• AES continues drawing up the chambers throughout the arcs. The task is about 60% complete. Several mismatches in the data we supplied them for chamber position resulted in a need to reestablish the baseline.
• We delivered data on positions of stands and cartridges to the Alignment Group for use in their planned network expansion and survey
• The Alignment Group established their expanded network in the FEL Enclosure. They will lay down the position of stands during short intervals over the next month.
• We started design of the Arc Dipole Stands

WBS 10 (Wiggler):
Final commissioning data was taken on the second wiggler of the optical klystron this month. This data gives us the data we need for control scripts for the wiggler power supply. The data is being analyzed. We hope to start commissioning data collection on the dispersion section next month.

We also finished analyzing the commissioning data for the second wiggler. The corrector field seems to be a bit stronger than it should be, leading to a 3 G DC component in the field. The initial commissioning values of the corrector fields will be corrected to account for this. We are now taking data on the vertical field to put a lower limit on the DC remnant in that direction.

WBS 11 (Optics):
The highlight of the month was a test of the 3" HR deformable mirror assembly (DMA) in the mirror test stand (MTS). Initial experiments were done without using a laser to induce a thermal deformation on the surface. More experiments were done where we loaded the front surface with the output from a multimode Nd:YAG laser. To date, analyses to date reveal the following:

With or without surface heating, the radius of curvature (ROC) can be changed linearly by varying just one parameter. This will simplify our feedback control.

Our technique for inducing the change in the radius of curvature does not create any additional aberrations to the mirror surface. There is no indication of "print-through" of the cooling channels on the mirror surface.

With surface heating of the mirror at levels above that expected in the IR Upgrade, we can control the ROC over the range we require (several meters change from a starting ROC of 17.5m).

Still under investigation is to what degree the ROC control during surface heating mitigates the thermally-induced spherical aberration, particularly as a function of input heat flux. This is to further benchmark the predictions from our earlier work, as well as from the FEA modeling done last year by AES.

We are very pleased with the results to date, and wish to acknowledge the efforts of AES and our Mechanical Engineering group in designing and fabricating the DMA, and to the I&C group as well for their assistance in terminating the instrument and heater cabling.

We held a teleconference with a potential vendor for the high power (50kW) laser beam dump, as well as for the backplane cooled insertable mirrors, we intend to send them some of our conceptual drawings. Design of the insertable mirror assemblies continues. We received a budgetary estimate on the leadscrews for the optical cavity assemblies, and plan to procure them by month’s end.

We continue to procure parts to build a pulse compressor for the "waste" drive laser IR output.

Based on experiments done at one vendor's location, we have requested a budgetary estimate for the oscillator and timing electronics. This laser would initially be tested at 74.85 MHz, then reconfigured to operate at 748.5 MHz.

After receiving budgetary estimates from the leadscrew vendor, we have modified the drawings; they were redone and sent to procurement . We are also pressing forward with plans to finish specifications for the outcouplers and get them procured. Design of the insertable mirror assemblies continues. We have continued work on the OCMMS, and have begun looking into low-cost ways to get the OCMMS HeNe beam size into EPICS. Fortunately, it appears we can use solutions developed outside the lab to do this.

We have submitted a requisition for the ultrafast laser that will be used to set the cavity length and perform other FEL diagnostics.

We held a meeting on the overall design of the optical transport system. Some excellent observations were made and are being incorporated into specific subassemblies. On the whole, the team agrees with the system plan. We worked on the overall design and installation schedule for integration with the rest of the machine.

We had a service call for our laser interferometer. The source laser was replaced and the imaging camera was realigned. We are testing the hardware as this report was written. We received actively stabilized vibration isolation legs for the mirror test stand optical table. We plan to install then today or early next week. We determined the flow rate through the 3" deformable mirror assembly to be 0.3 gal/min, quite modest.

We have received a sample of the fiber we will use as part of a pulse compressor, and optics have been set up before the drive laser to determine the damage threshold.

The drawings for the optical cavity assemblies were released for procurement this month. We continue work on the insertable mirror drawings. We are awaiting quotes on some of the mirrors for the OCMMS before final release of the drawings. The other long lead item (the calcium fluoride viewports) were ordered. We received the pillow blocks, another major component for the optical cavity, this week. These will be cleaned for UHV service next week. We received the modified ultraviewer components; and discovered the brackets had been bead blasted. After discussions with various vacuum-savvy folks, we have decided to have them remade.

We held more meetings on collimator designs, LCW requirements in the vault and upstairs, and electrical schematics for the optical cavity assemblies.

The pump laser for the mirror test stand (MTS) and the optical table supporting the laser interferometer and the MTS have improved vibration-isolation. This, and additional stiffening of the MTS floor have greatly lowered the amount of noise we see on the interferograms. Checks of the laser interferometer after the service done last week indicate that it is fixed.

UV FEL activities
We have the first estimates for the 40 K helium flow required to cool the sapphire outcoupler. Flow is quite modest, ~ 1 gm/sec, even in the case of rather highly absorbing (1000 ppm) mirrors. Considerable work remains, but this first estimate is encouraging.

UV FEL Microengineering Tool Development (Aerospace Corporation subcontract)
During this reporting period, work on the JLAB microengineering project proceeded with the continued design of the experimental station and the initiation of a small task, that explores a new laser material processing approach where ion implantation technology and a frequency up-converted IR FEL (e.g. 532 nm) are combined for mutual benefit. The merged processing approach permits the fabrication of patterned micro/nanometer scale structures, in semiconductors, over large areas (>400 mm dia). This task is in the exploratory research phase and, if proven feasible, could be easily transferred to the FEL as a viable light driven process that capitalizes on the nanometer scale depth control available in ion implantation. The merged material processing approach strongly benefits from the large average power of the FEL and that the wavelength can be tuned.

The design of the experimental station is currently in the phase where capability trade-offs are being addressed for the following critical subsystems. Optical delivery and control, piece-part motion control, vision and CADCAM. The team has been conducting literature and vendor research along with some modeling to arrive at a viable set of machine capabilities that can be translated to actual specifications. The design process started with a general wish-list of capabilities, that was refined to identify incompatibilities and then further delineated to address the trade-offs. We anticipate that these trade-off studies to be complete by the end of March with the goal of arriving at a system design in the June time frame.

An Aerospace internal Project Initiation Management Review (PIMR) was conducted for the benefit of upper management and the contract office. The PIMR addressed issues of this project regarding, deliverable, schedule and the impact to local staffing and the need and use of in-house facilities. Recommendations from this review will be implemented to strengthen the overall project and the approach.